# Kandinsky Conformal Prediction:

### v1

@@ 3.1 Applications
$\mathbb{P}_{(X_{n+1},Y_{n+1})\sim{\cal D}_{T}}[Y_{n+1}\in{\cal C}(X_{n+1})]=1-\alpha,$
@@
Let $\alpha$ and $\delta$ be parameters in $(0,1)$, and ${\cal W}=\{\Phi(\cdot)^{T}\beta:\beta\in\mathbb{R}^{d}\}$ denote a class of linear weight functions over a bounded basis $\Phi:{\cal X}\times{\cal Y}\to\mathbb{R}^{d}$. Under the same assumptions on $\{(X_{i},Y_{i})\}_{i\in[n+1]}$, $\widetilde{S}$, and $\Phi$ in Theorem 3.1, there exists an absolute constant $C$ such that, with probability at least $1-\delta$ over the randomness of the calibration dataset $\{(X_{i},Y_{i})\}_{i\in[n]}$ and the noise $\{\varepsilon_{i}\}_{i\in[n]}$, the (randomized) prediction set ${\cal C}$, given by Algorithms 1 and 2, satisfies
$\left|\mathbb{P}_{(X_{n+1},Y_{n+1})\sim{\cal D}_{T},\varepsilon_{n+1}}[Y_{n+1}\in{\cal C}(X_{n+1})]-(1-\alpha)\right|\leq B\left(C\sqrt{\frac{d}{n}}+\frac{d}{n}+\max\{\alpha,1-\alpha\}\sqrt{\frac{2\ln(4/\delta)}{n}}\right),$
where $(X_{n+1},Y_{n+1})$ are drawn independently from a distribution ${\cal D}_{\text{T}}$, for every distribution ${\cal D}_{\text{T}}$ such that $\frac{d\mathbb{P}_{{\cal D}_{T}}}{d\mathbb{P}_{{\cal D}}}\in{\cal W}$ and $\left|\frac{d\mathbb{P}_{{\cal D}_{T}}}{d\mathbb{P}_{{\cal D}}}(x,y)\right|\leq B$ for any $x\in{\cal X}$ and $y\in{\cal Y}$.
@@ 4 Test-time Quantile Regression
${\cal C}(X_{n+1})=\{y:{\widetilde{S}}(X_{n+1},y,\varepsilon_{n+1})\leq\widehat{q}_{y}(X_{n+1},y)\},$
@@
Let $\alpha$ be a parameter in $(0,1)$, and let ${\cal W}=\{\Phi(\cdot)^{T}\beta:\beta\in\mathbb{R}^{d}\}$ denote a class of linear weight functions over a basis $\Phi:{\cal X}\times{\cal Y}\to\mathbb{R}^{d}$. Assume that the data $\{(X_{i},Y_{i})\}_{i\in[n+1]}$ are drawn i.i.d. from a distribution ${\cal D}$, $\{\varepsilon_{i}\}_{i\in[n+1]}$ are drawn i.i.d. from a distribution ${\cal D}_{\text{rn}}$, independently from the dataset, and the distribution of ${\widetilde{S}}\left(X,Y,\varepsilon\right)\mid\Phi(X,Y)$ is continuous. Then, for any $w\in{\cal W}$, the prediction set given by Algorithm 3 satisfies
@@ 5 Experiments
We empirically evaluate the conditional coverage of Kandinsky conformal prediction on real-world tasks with natural groups: income prediction across US states (Ding et al., 2021) and toxic comment detection across demographic groups (Borkan et al., 2019; Koh et al., 2021). The data is divided into a training set for learning the base predictor, a calibration set for learning the conformal predictor, and a test set for evaluation. We repeat all experiments 100 times with reshuffled calibration and test sets.
@@
$\operatorname{M}(G,{\cal C})\coloneqq\frac{\sum_{i=n+1}^{n+m}\mathbbm{1}\left\{Y_{i}\not\in{\cal C}(X_{i})\right\}\cdot\mathbbm{1}\left\{Z_{i}\in G\right\}}{\sum_{i=n+1}^{n+m}\mathbbm{1}\left\{Z_{i}\in G\right\}}.$
@@ 5.2 CivilComments: A Classification Example
Kandinsky conformal prediction achieves the best group average coverage deviation when the calibration set contains an average of over 500 samples per group, as shown in Figure 3(d). Class-conditional conformal prediction achieves the best performance with 250 samples per group. However, unlike our approach, its coverage deviation does not improve with larger sample sizes. This is because Kandinsky considers the full set of groups, which are more fine-grained than classes. In Figures 3(e) and 3(f), we examine conditional miscoverage and prediction set sizes for each group under the setting of 4000 group average samples. Overall, Kandinsky generates prediction sets with conditional miscoverage closest to the target level for most groups while maintaining moderate sizes. We achieve the best calibrated coverage for groups with more than 2000 samples.
(e) CivilComments. Conditional miscoverage for each group. Target miscoverage is 0.1. Error bars are standard deviations of 100 runs.
(f) CivilComments. Average prediction set sizes for each group. Error bars are standard deviations of 100 runs.
Figure 3: Results. (a)(b)(c) ACSIncome. We vary the number of groups in calibration and test sets while controlling samples per group. (d) CivilComments. We redistribute samples between calibration and test sets to vary the number of calibration samples. (e)(f) CivilComments. There are 4000 calibration samples per group on average. We annotate the calibration sample size for each group. Groups with positive labels (toxic) are represented by y:1.
@@ A.1 Proofs from Section 3
Theorem 3.1 (Restated).
Let $\alpha$ and $\delta$ be parameters in $(0,1)$, and ${\cal W}=\{\Phi(\cdot)^{T}\beta:\beta\in\mathbb{R}^{d}\}$ denote a class of linear weight functions over a bounded basis $\Phi:{\cal X}\times{\cal Y}\to\mathbb{R}^{d}$. Assume that the data $\{(X_{i},Y_{i})\}_{i\in[n+1]}$ are drawn i.i.d. from a distribution ${\cal D}$, $\{\varepsilon_{i}\}_{i\in[n+1]}$ are drawn i.i.d. from a distribution ${\cal D}_{\text{rn}}$, independently from the dataset, and the distribution of ${\widetilde{S}}\left(X,Y,\varepsilon\right)\mid\Phi(X,Y)$ is continuous. Then, there exists an absolute constant $C$ such that, with probability at least $1-\delta$ over the randomness of the calibration dataset $\{(X_{i},Y_{i})\}_{i\in[n]}$ and the noise $\{\varepsilon_{i}\}_{i\in[n]}$, the (randomized) prediction set ${\cal C}$, given by Algorithms 1 and 2, satisfies
$\left|\mathbb{E}_{(X_{n+1},Y_{n+1})\sim{\cal D},\varepsilon_{n+1}\sim{\cal D}_{\text{rn}}}\left[w(X_{n+1},Y_{n+1})\left(\mathbbm{1}\left\{Y_{n+1}\in{\cal C}(X_{n+1})\right\}-(1-\alpha)\right)\right]\right|\leq$
${}\|w\|_{\infty}\left(C\sqrt{\frac{d}{n}}+\frac{d}{n}+\max\{\alpha,1-\alpha\}\sqrt{\frac{2\ln(4/\delta)}{n}}\right),$
for every $w\in{\cal W}$.
@@
Denote ${\widetilde{S}}\left(X_{i},Y_{i},\varepsilon_{i}\right)$ by ${\widetilde{S}}_{i}$. Then $\{(X_{i},Y_{i},{\widetilde{S}}_{i})\}_{i\in[n+1]}$ are independent and identically distributed.
@@
$0\in~{}$ $\partial L(\eta)\big{|}_{\eta=0}=\left\{\frac{1}{n}\sum_{i=1}^{n}w(X_{i},Y_{i})\gamma_{i}~{}\bigg{|}~{}\gamma_{i}=\begin{cases}\alpha-\mathbbm{1}\{{\widetilde{S}}_{i}>\widehat{q}(X_{i},Y_{i})\},&\widehat{q}(X_{i},Y_{i})\neq{\widetilde{S}}_{i},\\ \alpha_{i}\in[\alpha-1,\alpha],&\widehat{q}(X_{i},Y_{i})=\widetilde{S}_{i}\end{cases}\right\}.$
@@
$\mathbb{E}\left[h(X_{n+1},Y_{n+1},{\widetilde{S}}_{n+1})\right]-\frac{1}{n}\sum_{i=1}^{n}h(X_{i},Y_{i},{\widetilde{S}}_{i})$ $\leq 2\mathbb{E}_{D}\left[{\cal R}({\cal H}_{2}\circ D)\right]+\max\left\{\alpha,1-\alpha\right\}\|w\|_{\infty}\sqrt{\frac{2\ln(2/\delta)}{n}}$
@@
$\frac{1}{n}\sum_{i=1}^{n}h(X_{i},Y_{i},{\widetilde{S}}_{i})-\mathbb{E}\left[h(X_{n+1},Y_{n+1},{\widetilde{S}}_{n+1})\right]\leq 2C^{\prime}\|w\|_{\infty}\sqrt{\frac{2d}{n}}+\max\left\{\alpha,1-\alpha\right\}\|w\|_{\infty}\sqrt{\frac{2\ln(2/\delta)}{n}}.$
@@
$\left|\mathbb{E}\left[h(X_{n+1},Y_{n+1},{\widetilde{S}}_{n+1})\right]-\frac{1}{n}\sum_{i=1}^{n}h(X_{i},Y_{i},{\widetilde{S}}_{i})\right|\leq 2C^{\prime}\|w\|_{\infty}\sqrt{\frac{2d}{n}}+\max\left\{\alpha,1-\alpha\right\}\|w\|_{\infty}\sqrt{\frac{2\ln(4/\delta)}{n}}.$
@@
$\left|\mathbb{E}\left[w(X_{n+1},Y_{n+1})\left(\mathbbm{1}\left\{{\widetilde{S}}_{n+1}\leq q(X_{n+1},Y_{n+1})\right\}-(1-\alpha)\right)\right]-\frac{1}{n}\sum_{i=1}^{n}w(X_{i},Y_{i})\left(\alpha-\mathbbm{1}\left\{q(X_{i},Y_{i})-{\widetilde{S}}_{i}<0\right\}\right)\right|=$
$\left|\mathbb{E}\left[w(X_{n+1},Y_{n+1})\left(\alpha-\mathbbm{1}\left\{q(X_{n+1},Y_{n+1})-{\widetilde{S}}_{n+1}<0\right\}\right)\right]-\frac{1}{n}\sum_{i=1}^{n}w(X_{i},Y_{i})\left(\alpha-\mathbbm{1}\left\{q(X_{i},Y_{i})-{\widetilde{S}}_{i}<0\right\}\right)\right|\leq$
@@
$\begin{split}&\left|\mathbb{E}_{(X_{n+1},Y_{n+1})\sim{\cal D},\varepsilon_{n+1}\sim{\cal D}_{\text{rn}}}\left[w(X_{n+1},Y_{n+1})\left(\mathbbm{1}\left\{{\widetilde{S}}_{n+1}\leq\widehat{q}(X_{n+1},Y_{n+1})\right\}-(1-\alpha)\right)\right]\right|\leq\\ &\left|\frac{1}{n}\sum_{i=1}^{n}w(X_{i},Y_{i})\left(\alpha-\mathbbm{1}\left\{\widehat{q}(X_{i},Y_{i})-{\widetilde{S}}_{i}<0\right\}\right)\right|+2C^{\prime}\|w\|_{\infty}\sqrt{\frac{2d}{n}}+\max\left\{\alpha,1-\alpha\right\}\|w\|_{\infty}\sqrt{\frac{2\ln(4/\delta)}{n}}\leq\\ &\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}\left\{\widehat{q}(X_{i}.Y_{i})={\widetilde{S}}_{i}\right\}|w(X_{i},Y_{i})|+2C^{\prime}\|w\|_{\infty}\sqrt{\frac{2d}{n}}+\max\left\{\alpha,1-\alpha\right\}\|w\|_{\infty}\sqrt{\frac{2\ln(4/\delta)}{n}}\leq\\ &\frac{1}{n}\|w\|_{\infty}\sum_{i=1}^{n}\mathbbm{1}\left\{\widehat{q}(X_{i}.Y_{i})={\widetilde{S}}_{i}\right\}+2C^{\prime}\|w\|_{\infty}\sqrt{\frac{2d}{n}}+\max\left\{\alpha,1-\alpha\right\}\|w\|_{\infty}\sqrt{\frac{2\ln(4/\delta)}{n}}.\end{split}$ (5)
@@
$0$
The last equation holds because the distribution of $\left({\widetilde{S}}_{j_{1}},...,{\widetilde{S}}_{j_{d+1}}\right)\Bigm{|}\Phi(X_{1},Y_{1}),...,\Phi(X_{n},Y_{n})$ is continuous, and $\operatorname{RowSpace}\left(\left[\Phi(X_{j_{1}},Y_{j_{1}}),...,\Phi(X_{j_{d+1}},Y_{j_{d+1}})\right]\right)$ is at most $d-$dimensional while $\left({\widetilde{S}}_{j_{1}},...,{\widetilde{S}}_{j_{d+1}}\right)$ is a $(d+1)-$dimensional vector. Therefore, with probability 1,
@@
$\left|\mathbb{E}_{(X_{n+1},Y_{n+1})\sim{\cal D},\varepsilon_{n+1}\sim{\cal D}_{\text{rn}}}\left[w(X_{n+1},Y_{n+1})\left(\mathbbm{1}\left\{{\widetilde{S}}_{n+1}\leq\widehat{q}(X_{n+1},Y_{n+1})\right\}-(1-\alpha)\right)\right]\right|\leq$

### v2

@@ 3.1 Applications
$\mathbb{P}_{{\cal D}_{T}}[Y_{n+1}\in{\cal C}(X_{n+1};\varepsilon_{n+1})\mid\{X_{i},Y_{i},\varepsilon_{i}\}_{i\in[n]}]\approx 1-\alpha,$
@@
Let $\alpha$ and $\delta$ be parameters in (0,1)(0,1)( 0 , 1 ), and ${\cal W}=\{\Phi(\cdot)^{T}\beta:\beta\in\mathbb{R}^{d}\}$ denote a class of linear weight functions over a bounded basis $\Phi:{\cal X}\times{\cal Y}\to\mathbb{R}^{d}$. Under the same assumptions on $\{(X_{i},Y_{i},\varepsilon_{i})\}_{i\in[n]}$, $\widetilde{S}$, and $\Phi$ in Theorem 3.1, there exists an absolute constant $C$ such that, for every distribution ${\cal D}_{\text{T}}$ such that $\frac{d\mathbb{P}_{{\cal D}_{T}}}{d\mathbb{P}_{{\cal D}}}\in{\cal W}$ and $\left|\frac{d\mathbb{P}_{{\cal D}_{T}}}{d\mathbb{P}_{{\cal D}}}(x,y)\right|\leq B$ for any $x\in{\cal X}$ and $y\in{\cal Y}$, with probability at least $1-\delta$ over the randomness of the calibration dataset $\{(X_{i},Y_{i})\}_{i\in[n]}$ and the noise $\{\varepsilon_{i}\}_{i\in[n]}$, the (randomized) prediction set ${\cal C}$, given by Algorithms 1 and 2, satisfies
$|\mathbb{P}[Y_{n+1}\in{\cal C}(X_{n+1};\varepsilon_{n+1})\mid\{X_{i},Y_{i},\varepsilon_{i}\}_{i\in[n]}]-(1-\alpha)|\leq B\left(C\sqrt{\frac{d}{n}}+\frac{d}{n}+\max\{\alpha,1-\alpha\}\sqrt{\frac{2\ln(4/\delta)}{n}}\right),$
where $(X_{n+1},Y_{n+1})$ are drawn independently from the distribution ${\cal D}_{\text{T}}$, and $\varepsilon_{n+1}$ are drawn independently from the distribution ${\cal D}_{\text{rn}}$.
@@ 4 Test-time Quantile Regression
${\cal C}(X_{n+1};\varepsilon_{n+1})=\{y:{\widetilde{S}}(X_{n+1},y,\varepsilon_{n+1})\leq\widehat{q}_{y}(X_{n+1},y)\},$
@@
Let $\alpha$ be a parameter in (0,1)(0,1)( 0 , 1 ), and let ${\cal W}=\{\Phi(\cdot)^{T}\beta:\beta\in\mathbb{R}^{d}\}$ denote a class of linear weight functions over a basis $\Phi:{\cal X}\times{\cal Y}\to\mathbb{R}^{d}$. Assume that the data $\{(X_{i},Y_{i})\}_{i\in[n+1]}$ are drawn i.i.d. from a distribution ${\cal D}$, $\{\varepsilon_{i}\}_{i\in[n+1]}$ are drawn i.i.d. from a distribution ${\cal D}_{\text{rn}}$, independently from the dataset, and the distribution of ${\widetilde{S}}\left(X,Y,\varepsilon\right)\mid\Phi(X,Y)$ is continuous. Then, for any $w\in{\cal W}$, the prediction set given by Algorithm 3 satisfies
@@ 5 Experiments
Figure 3: Main results. (a)(b)(c) ACSIncome. We vary the number of groups in calibration and test sets while controlling samples per group. (d) CivilComments. We redistribute samples between calibration and test sets to vary the number of calibration samples.
(a) CivilComments. Conditional miscoverage for each group. Target miscoverage is 0.1. Error bars are standard deviations of 100 runs.
(b) CivilComments. Average prediction set sizes for each group. Error bars are standard deviations of 100 runs.
Figure 4: Group-wise metrics on CivilComments. There are 4000 calibration samples per group on average. We annotate the calibration sample size for each group. Groups with positive labels (toxic) are represented by y:1.
(a) The distribution of group average coverage deviation across 10 prompts.
(b) The distribution of average prediction set sizes across 10 prompts.
Figure 5: MCQA. We experiment with the prediction of the LLaMA-13B model across 10 one-shot prompts.
We empirically evaluate the conditional coverage of Kandinsky conformal prediction on real-world tasks with natural groups: income prediction across US states (Ding et al., 2021), toxic comment detection across demographic groups (Borkan et al., 2019; Koh et al., 2021), and multiple-choice question answering across subjects (Hendrycks et al., 2021; Kumar et al., 2023). For the first two datasets, the data is divided into a training set for learning the base predictor, a calibration set for learning the conformal predictor, and a test set for evaluation. We repeat all experiments 100 times with reshuffled calibration and test sets. While for multile-choice question answering, we adopt LLaMA-13B (Touvron et al., 2023) as a blackbox base predictor.
@@
$\operatorname{M}(G,{\cal C})\coloneqq\frac{\sum_{i=n+1}^{n+m}\mathbbm{1}\left\{Y_{i}\not\in{\cal C}(X_{i},\varepsilon_{i})\right\}\cdot\mathbbm{1}\left\{Z_{i}\in G\right\}}{\sum_{i=n+1}^{n+m}\mathbbm{1}\left\{Z_{i}\in G\right\}}.$
@@ 5.2 CivilComments: A Classification Example
Kandinsky conformal prediction achieves the best group average coverage deviation when the calibration set contains an average of over 500 samples per group, as shown in Figure 3(d). Class-conditional conformal prediction achieves the best performance with 250 samples per group. However, unlike our approach, its coverage deviation does not improve with larger sample sizes. This is because Kandinsky considers the full set of groups, which are more fine-grained than classes. In Figures 4(a) and 4(b), we examine conditional miscoverage and prediction set sizes for each group under the setting of 4000 group average samples. Overall, Kandinsky generates prediction sets with conditional miscoverage closest to the target level for most groups while maintaining moderate sizes. We achieve the best calibrated coverage for groups with more than 2000 samples.
5.3 Multi-Choice Question Answering: A Blackbox Example
We adopt a setup where the model is a blackbox. We have access to the prediction of the LLaMA-13B (Touvron et al., 2023) model for the MMLU (Hendrycks et al., 2021) benchmark, a multiple-choice question answering (MCQA) task. Kumar et al. (2023) take a subset of MMLU, prompt the model with one-shot examples, and provide the predicted probabilities for four choices. Ten different prompts are used for the same question, producing ten parallel predictions. We take as input the output probabilities of the model as logged by Kumar et al. (2023), while assuming unavailability of both the model and the original multiple-choice questions. Therefore, Kandinsky conformal prediction (FY) is applicable to the task while Kandinsky (XY) is not. Kumar et al. (2023) group the questions into 16 subjects, spanning the fields of computer science, business, and medicine. We further classify the subjects into two categories according to their difficulty, which is measured by the accuracy of the LLaMA model on those subjects as reported by Kumar et al. (2023). The basic group contains subjects with accuracy over 40% while the remaining subjects are in the advanced group. The details are specified in Section C.3. The model’s confidence calibration differs between the two groups. We use the APS score function for conformal prediction. We repeat the experiment for each of the ten prompts.
Compared with split conformal prediction, Kandinsky conformal prediction achieves the desired coverage with lower deviation error per group in average, as shown in Figure 5(a). Figure 5(b) shows it also produces less redundant prediction sets. While conservative conformal prediction has comparable group average coverage deviation with Kandinsky, it is achieved by overly conservative prediction sets with significantly larger sizes.
@@ A.1 Proofs from Section 3
We first show in the following Lemma that for a fixed $w\in{\cal W}$, the expected weighted coverage deviation is bounded with high probability. Note that this statement provides a bound for a single weight function.
Lemma A.1.
Let $\alpha$ and $\delta$ be parameters in (0,1)(0,1)( 0 , 1 ), and ${\cal W}=\{\Phi(\cdot)^{T}\beta:\beta\in\mathbb{R}^{d}\}$ denote a class of linear weight functions over a bounded basis $\Phi:{\cal X}\times{\cal Y}\to\mathbb{R}^{d}$. Assume that the data $\{(X_{i},Y_{i})\}_{i\in[n+1]}$ are drawn i.i.d. from a distribution ${\cal D}$, $\{\varepsilon_{i}\}_{i\in[n+1]}$ are drawn i.i.d. from a distribution ${\cal D}_{\text{rn}}$, independently from the dataset, and the distribution of ${\widetilde{S}}\left(X,Y,\varepsilon\right)\mid\Phi(X,Y)$ is continuous. Then, there exists an absolute constant $C$ such that, for every $w\in{\cal W}$, with probability at least $1-\delta$ over the randomness of the calibration dataset $\{(X_{i},Y_{i})\}_{i\in[n]}$ and the noise $\{\varepsilon_{i}\}_{i\in[n]}$, the (randomized) prediction set ${\cal C}$, given by Algorithms 1 and 2, satisfies
$\left|\mathbb{E}_{(X,Y)\sim{\cal D},\varepsilon\sim{\cal D}_{\text{rn}}}\left[w(X,Y)\left(\mathbbm{1}\left\{Y\in{\cal C}(X;\varepsilon)\right\}-(1-\alpha)\right)\right]\right|\leq\|w\|_{\infty}\left(C\sqrt{\frac{d}{n}}+\frac{d}{n}+\max\{\alpha,1-\alpha\}\sqrt{\frac{2\ln(4/\delta)}{n}}\right).$
@@
Denote ${\widetilde{S}}\left(X_{i},Y_{i},\varepsilon_{i}\right)$ by ${\widetilde{S}}_{i}$. Then $\{(X_{i},Y_{i},{\widetilde{S}}_{i})\}_{i\in[n]}$ are independent and identically distributed.
@@
0 0$\in$~0 ∈ $\partial L(\eta)\big{|}_{\eta=0}=\left\{\frac{1}{n}\sum_{i=1}^{n}w(X_{i},Y_{i})\gamma_{i}~\bigg{|}~\gamma_{i}=\begin{cases}\alpha-\mathbbm{1}\{{\widetilde{S}}_{i}>\widehat{q}(X_{i},Y_{i})\},&\widehat{q}(X_{i},Y_{i})\neq{\widetilde{S}}_{i},\\ \alpha_{i}\in[\alpha-1,\alpha],&\widehat{q}(X_{i},Y_{i})=\widetilde{S}_{i}\end{cases}\right\}.$
@@
$\mathbb{E}\left[h(X,Y,{\widetilde{S}})\right]-\frac{1}{n}\sum_{i=1}^{n}h(X_{i},Y_{i},{\widetilde{S}}_{i})$ $\leq 2\mathbb{E}_{D}\left[{\cal R}({\cal H}_{2}\circ D)\right]+\max\left\{\alpha,1-\alpha\right\}\|w\|_{\infty}\sqrt{\frac{2\ln(2/\delta)}{n}}$
@@
$\frac{1}{n}\sum_{i=1}^{n}h(X_{i},Y_{i},{\widetilde{S}}_{i})-\mathbb{E}\left[h(X,Y,{\widetilde{S}})\right]\leq 2C^{\prime}\|w\|_{\infty}\sqrt{\frac{2d}{n}}+\max\left\{\alpha,1-\alpha\right\}\|w\|_{\infty}\sqrt{\frac{2\ln(2/\delta)}{n}}.$
@@
$\left|\mathbb{E}\left[h(X,Y,{\widetilde{S}})\right]-\frac{1}{n}\sum_{i=1}^{n}h(X_{i},Y_{i},{\widetilde{S}}_{i})\right|\leq 2C^{\prime}\|w\|_{\infty}\sqrt{\frac{2d}{n}}+\max\left\{\alpha,1-\alpha\right\}\|w\|_{\infty}\sqrt{\frac{2\ln(4/\delta)}{n}}.$
@@
$\left|\mathbb{E}\left[w(X,Y)\left(\mathbbm{1}\left\{{\widetilde{S}}\leq q(X,Y)\right\}-(1-\alpha)\right)\right]-\frac{1}{n}\sum_{i=1}^{n}w(X_{i},Y_{i})\left(\alpha-\mathbbm{1}\left\{q(X_{i},Y_{i})-{\widetilde{S}}_{i}<0\right\}\right)\right|=$
$\left|\mathbb{E}\left[w(X,Y)\left(\alpha-\mathbbm{1}\left\{q(X,Y)-{\widetilde{S}}<0\right\}\right)\right]-\frac{1}{n}\sum_{i=1}^{n}w(X_{i},Y_{i})\left(\alpha-\mathbbm{1}\left\{q(X_{i},Y_{i})-{\widetilde{S}}_{i}<0\right\}\right)\right|\leq$
@@
$\begin{split}&\left|\mathbb{E}_{(X,Y)\sim{\cal D},\varepsilon\sim{\cal D}_{\text{rn}}}\left[w(X,Y)\left(\mathbbm{1}\left\{{\widetilde{S}}\leq\widehat{q}(X,Y)\right\}-(1-\alpha)\right)\right]\right|\leq\\ &\left|\frac{1}{n}\sum_{i=1}^{n}w(X_{i},Y_{i})\left(\alpha-\mathbbm{1}\left\{\widehat{q}(X_{i},Y_{i})-{\widetilde{S}}_{i}<0\right\}\right)\right|+2C^{\prime}\|w\|_{\infty}\sqrt{\frac{2d}{n}}+\max\left\{\alpha,1-\alpha\right\}\|w\|_{\infty}\sqrt{\frac{2\ln(4/\delta)}{n}}\leq\\ &\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}\left\{\widehat{q}(X_{i}.Y_{i})={\widetilde{S}}_{i}\right\}|w(X_{i},Y_{i})|+2C^{\prime}\|w\|_{\infty}\sqrt{\frac{2d}{n}}+\max\left\{\alpha,1-\alpha\right\}\|w\|_{\infty}\sqrt{\frac{2\ln(4/\delta)}{n}}\leq\\ &\frac{1}{n}\|w\|_{\infty}\sum_{i=1}^{n}\mathbbm{1}\left\{\widehat{q}(X_{i}.Y_{i})={\widetilde{S}}_{i}\right\}+2C^{\prime}\|w\|_{\infty}\sqrt{\frac{2d}{n}}+\max\left\{\alpha,1-\alpha\right\}\|w\|_{\infty}\sqrt{\frac{2\ln(4/\delta)}{n}}.\end{split}$ (6)
@@
0 0.0 .
[…]
@@
$\left|\mathbb{E}_{(X,Y)\sim{\cal D},\varepsilon\sim{\cal D}_{\text{rn}}}\left[w(X,Y)\left(\mathbbm{1}\left\{{\widetilde{S}}\leq\widehat{q}(X,Y)\right\}-(1-\alpha)\right)\right]\right|\leq$
@@
Using Lemma A.1 and a union bound, we prove a high probability bound on the expected coverage deviation for all weight functions in ${\cal W}$.
Theorem 3.1 (Restated).
Let $\alpha$ and $\delta$ be parameters in (0,1)(0,1)( 0 , 1 ), and ${\cal W}=\{\Phi(\cdot)^{T}\beta:\beta\in\mathbb{R}^{d}\}$ denote a class of linear weight functions over a bounded basis $\Phi:{\cal X}\times{\cal Y}\to\mathbb{R}^{d}$. Without loss of generality, assume $\|\Phi(x,y)\|_{\infty}\leq 1$ for all $x,y$. Assume that the data $\{(X_{i},Y_{i})\}_{i\in[n]}$ are drawn i.i.d. from a distribution ${\cal D}$, $\{\varepsilon_{i}\}_{i\in[n]}$ are drawn i.i.d. from a distribution ${\cal D}_{\text{rn}}$, independently from the dataset, and the distribution of ${\widetilde{S}}\left(X,Y,\varepsilon\right)\mid\Phi(X,Y)$ is continuous. Then, there exists an absolute constant $C$ such that, with probability at least $1-\delta$ over the randomness of the calibration dataset $\{(X_{i},Y_{i})\}_{i\in[n]}$ and the noise $\{\varepsilon_{i}\}_{i\in[n]}$, the (randomized) prediction set ${\cal C}$, given by Algorithms 1 and 2, satisfies
$\left|\mathbb{E}_{(X,Y)\sim{\cal D},\varepsilon\sim{\cal D}_{\text{rn}}}\left[w_{\beta}(X,Y)\left(\mathbbm{1}\left\{Y\in{\cal C}(X;\varepsilon)\right\}-(1-\alpha)\right)\right]\right|\leq\|\beta\|_{1}\left(C\sqrt{\frac{d}{n}}+\frac{d}{n}+\max\{\alpha,1-\alpha\}\sqrt{\frac{2\ln(4d/\delta)}{n}}\right),$
for every $w_{\beta}=\Phi(\cdot)^{T}\beta$.
Proof.
Let $\Phi(\cdot)=(\phi_{1}(\cdot),\cdots,\phi_{d}(\cdot))^{T}$, where $\phi_{j}\in\mathbb{R}^{{\cal X}\times{\cal Y}}$ for $1\leq j\leq d$. According to Lemma A.1, for all $1\leq j\leq d$, with probability at least $1-\delta/d$,
$\left|\mathbb{E}_{(X,Y)\sim{\cal D},\varepsilon\sim{\cal D}_{\text{rn}}}\left[\phi_{j}(X,Y)\left(\mathbbm{1}\left\{Y\in{\cal C}(X;\varepsilon)\right\}-(1-\alpha)\right)\right]\right|\leq C\sqrt{\frac{d}{n}}+\frac{d}{n}+\max\{\alpha,1-\alpha\}\sqrt{\frac{2\ln(4d/\delta)}{n}}.$
Then, we have the union bound. With probability at least $1-\delta$, for all $1\leq j\leq d$,
$\left|\mathbb{E}_{(X,Y)\sim{\cal D},\varepsilon\sim{\cal D}_{\text{rn}}}\left[\phi_{j}(X,Y)\left(\mathbbm{1}\left\{Y\in{\cal C}(X;\varepsilon)\right\}-(1-\alpha)\right)\right]\right|\leq C\sqrt{\frac{d}{n}}+\frac{d}{n}+\max\{\alpha,1-\alpha\}\sqrt{\frac{2\ln(4d/\delta)}{n}}.$
For any $w_{\beta}=\Phi(\cdot)^{T}\beta$,
$\left|\mathbb{E}_{(X,Y)\sim{\cal D},\varepsilon\sim{\cal D}_{\text{rn}}}\left[w_{\beta}(X,Y)\left(\mathbbm{1}\left\{Y\in{\cal C}(X;\varepsilon)\right\}-(1-\alpha)\right)\right]\right|\leq$
${}\sum_{j=1}^{d}\left|\mathbb{E}_{(X,Y)\sim{\cal D},\varepsilon\sim{\cal D}_{\text{rn}}}\left[\beta_{j}\phi_{j}(X,Y)\left(\mathbbm{1}\left\{Y\in{\cal C}(X;\varepsilon)\right\}-(1-\alpha)\right)\right]\right|\leq$
${}\sum_{j=1}^{d}\left|\beta_{j}\right|\cdot\left(C\sqrt{\frac{d}{n}}+\frac{d}{n}+\max\{\alpha,1-\alpha\}\sqrt{\frac{2\ln(4d/\delta)}{n}}\right)=$
${}\|\beta\|_{1}\left(C\sqrt{\frac{d}{n}}+\frac{d}{n}+\max\{\alpha,1-\alpha\}\sqrt{\frac{2\ln(4d/\delta)}{n}}\right).$